Show that if Zn = (a + b")
1
Expert's answer
2021-10-11T15:49:05-0400

if "z_n=(a^n+b^n)^{1\/n}" where "0<a<b" , then

"\displaystyle{\lim_{n\to \infin}}z_n=b"


"(a^n+b^n)^{1\/n}=b((a\/b)^n+1)^{1\/n}"


"\displaystyle{\lim_{n\to \infin}}(a/b)^n=0"


"\displaystyle{\lim_{n\to \infin}}((a/b)^n+1)^{1/n}=0"


"\displaystyle{\lim_{n\to \infin}}(b((a/b)^n+1)^{1/n})=b\cdot1=b"